# Negative Token Merging: Image-based Adversarial Feature Guidance

### v1

@@ Abstract
Text-based adversarial guidance using a negative prompt has emerged as a widely adopted approach to push the output features away from undesired concepts. While useful, performing adversarial guidance using text alone can be insufficient to capture complex visual concepts and avoid undesired visual elements like copyrighted characters. In this paper, for the first time we explore an alternate modality in this direction by performing adversarial guidance directly using visual features from a reference image or other images in a batch. In particular, we introduce negative token merging (NegToMe), a simple but effective training-free approach which performs adversarial guidance by selectively pushing apart matching semantic features (between reference and output generation) during the reverse diffusion process. When used w.r.t. other images in the same batch, we observe that NegToMe significantly increases output diversity (racial, gender, visual) without sacrificing output image quality. Similarly, when used w.r.t. a reference copyrighted asset, NegToMe helps reduce visual similarity with copyrighted content by 34.57%. NegToMe is simple to implement using just few-lines of code, uses only marginally higher ($<4\%$) inference times and generalizes to different diffusion architectures like Flux, which do not natively support the use of a separate negative prompt. Code is available at https://negtome.github.io.
(a) Adversarial Guidance across Different Outputs: State-of-the-art diffusion models are observed to suffer from limited diversity (e.g., ethnic, racial, gender etc.). NegToMe can be used to improve output diversity by adversarially guiding each image away from each other during reverse diffusion process.
@@ 1 Introduction
In this paper, we explore an alternate modality in this direction by performing adversarial guidance using images. Our key intuition is that even if describing the undesired concepts is not effective or feasible as text (child in park for Fig. 2), we can directly use the visual features from a reference image in order to adversarially guide the generation process. For instance in Fig. 2, instead of trying to exhaustively describe the child’s attire, placement, pose, background etc., we wish to directly use the visual features from the reference image to guide the generation process. Similarly, in cases where a negative prompt alone is not sufficient (e.g., copyrighted characters in Fig. 1), we can better guide the generation away from undesired concepts by directly using the character images for adversarial guidance.
To this end, we propose negative token merging (NegToMe), a simple and training-free approach which achieves adversarial guidance by selectively pushing apart semantically matching features (between the reference image and generated outputs) during the reverse diffusion process. NegToMe is easy to implement with only a few lines of code and can be integrated into various diffusion architectures, enabling a range of custom applications (see Fig. 2); 1) adversarial guidance for visually complex concepts (e.g., ‘child in a park’), 2) style-guidance for excluding specific artistic elements, 3) enhancing output aesthetics using a blurry reference image, and, 4) object feature interpolation or extrapolation e.g., between a kitten and young cat (Fig. 2) to inter-extrapolate visual features for cat age, size etc.
Figure 2: Image-based adversarial guidance. Simply adjusting the reference allows for a range of custom applications with NegToMe (e.g., improve output quality by using a blurry reference).
In particular, to demonstrate the practical usefulness of the proposed approach, we identify two prominent use-cases of NegToMe; 1) improving output diversity (when performed w.r.t other images in same batch), and, 2) reducing visual similarity to copyrighted characters (when performed w.r.t a copyrighted RAG database). For instance, it has been empirically shown [29] that state-of-the-art diffusion models often struggle from the problem of limited output diversity (e.g., limited racial, gender diversity for a prompt for a person in Fig. 1). The use of NegToMe across a batch, inherently helps addresses this problem by pushing visual features of each image away from the each other during reverse diffusion process. Through both qualitative and quantitative results (refer Sec. 4.1) we show that NegToMe helps significantly improve output diversity (racial, gender, ethnic, visual) without requiring any training or finetuning.
@@ 3 Negative Token Merging
Semantic Token Matching. A key idea behind NegToMe is to push each output token (source) away from its semantically closest token (target) in the reference image during the reverse diffusion process. This requires accurate computation of semantic token-to-token correspondences (e.g., head of child in Fig. 3) between the generated tokens and the tokens in the reference image. Luckily, we can leverage the rich semantic structure of the intermediate diffusion features [45, 52] itself to compute cross-image token-token similarities using noisy latent features itself.

### v2

@@ Abstract
Text-based adversarial guidance using a negative prompt has emerged as a widely adopted approach to steer diffusion models away from producing undesired concepts. While useful, performing adversarial guidance using text alone can be insufficient to capture complex visual concepts or avoid specific visual elements like copyrighted characters. In this paper, for the first time we explore an alternate modality in this direction by performing adversarial guidance directly using visual features from a reference image or other images in a batch. We introduce negative token merging (NegToMe), a simple but effective training-free approach which performs adversarial guidance through images by selectively pushing apart matching visual features between reference and generated images during the reverse diffusion process. By simply adjusting the used reference, NegToMe enables a diverse range of applications. Notably, when using other images in same batch as reference, we find that NegToMe significantly enhances output diversity (e.g., racial, gender, visual) by guiding features of each image away from others. Similarly, when used w.r.t. copyrighted reference images, NegToMe reduces visual similarity to copyrighted content by 34.57%. NegToMe is simple to implement using just few-lines of code, uses only marginally higher ($<4\%$) inference time and is compatible with different diffusion architectures, including those like Flux, which don’t natively support the use of a negative prompt. Code is available at https://negtome.github.io.
(a) Adversarial Guidance across Different Outputs: State-of-the-art diffusion models are observed to suffer from limited diversity (e.g., ethnic, racial, gender etc.). NegToMe can be used to improve output diversity by simply guiding the features of each image away from each other during reverse diffusion.
@@ 1 Introduction
To address this challenge, we propose negative token merging (NegToMe), a simple training-free approach that performs adversarial guidance directly using images instead of text. Our key insight is that even when text is not sufficient to capture visually complex concepts (child in park for Fig. 2), we can directly use the visual features from a reference image to adversarially guide the generation process. For instance in Fig. 2, instead of trying to exhaustively describe the child’s attire, placement, pose, background etc., we can directly use the reference image for better guidance. Similarly, in cases where a negative prompt alone is not sufficient (e.g., copyrighted characters in Fig. 1), we can better guide the generation away from undesired concepts by directly using the character images for adversarial guidance.
NegToMe is easy to implement and can be incorporated into various diffusion architectures using only a few lines of code (Alg. 1). By simply adjusting the used reference, NegToMe enables a range of custom applications (see Fig. 2); 1) adversarial guidance for visually complex concepts (e.g., ‘child in a park’), 2) style-guidance for excluding specific artistic elements, 3) enhancing output aesthetics by guiding away from a blurry reference, and, 4) object feature interpolation or extrapolation by guiding the outputs towards or away from the reference e.g., between a kitten and young cat (Fig. 2) to inter-extrapolate features for cat age, size etc.
Figure 2: Image-based adversarial guidance. NegToMe enables directly using images (instead of negative prompt alone) for adversarial guidance. By simply adjusting the used reference, NegToMe allows for a range of custom applications, 1) adversarial guidance for visually complex concepts to improve diversity, 2) Style control for excluding specific artistic styles (e.g., Van Gogh) from generated images, 3) improving output quality by guiding away from a blurry reference, 4) Object feature interpolation or extrapolation by guiding the outputs towards or away from the reference etc.
In particular, to demonstrate the practical usefulness of the proposed approach, we identify two prominent use-cases of NegToMe; 1) improving output diversity (when using other images in same batch as reference), and, 2) reducing visual similarity to copyrighted characters (when performed w.r.t a copyrighted RAG database). For instance, it has been empirically shown [29] that state-of-the-art diffusion models often struggle from the problem of limited output diversity (e.g., limited racial, gender diversity for a prompt for a person in Fig. 1). The use of NegToMe across a batch, inherently helps addresses this problem by pushing visual features of each image away from the each other during reverse diffusion process. Through both qualitative and quantitative results (refer Sec. 4.1) we show that NegToMe helps significantly improve output diversity (racial, gender, ethnic, visual) without requiring any training or finetuning.
@@ 3 Negative Token Merging
Semantic Token Matching. A key idea behind NegToMe is to push each output token (source) away from its semantically closest token (target) in the reference image during the reverse diffusion process. This requires computation of semantic token-to-token correspondences between the generated tokens and the tokens in the reference image. Luckily, we can leverage the rich semantic structure of intermediate diffusion features [45, 52] to compute cross-image token-token similarities using noisy features itself.